The sum of two numbers is 119. One number is 14 more than half another number. Find the numbers.
Answer by checkley77(12844)

X+Y=119 OR X=119-Y
X=14+Y/2
119-Y=14+Y/2
119-14=Y/2+Y
105=1.5Y
Y=1.05/1.5
Y=70 ANS.
X+70=119
X=119-70
X=49 ANS.
PROOF:
49=14+70/2
49=14+35
49=49